# Localization and entanglement characterization of edge states in HgTe quantum wells in a finite strip geometry

### v2

@@ Abstract
Quantum information measures are proposed to analyze the structure of near-gap electronic states in HgTe quantum wells in a strip geometry $(x,y)\in(-\infty,\infty)\times[0,L]$ of finite width $L$. This allows us to establish criteria for distinguishing edge from bulk states in the topological insulator phase, including the transition region and cutoff of the wave number $k_{x}$ where edge states degenerate with bulk states. Qualitative and quantitative information on the near-gap Hamiltonian eigenstates, obtained by tight-binding calculations, is extracted from localization measures, like the inverse participation ratio (IPR), and entanglement entropies of the reduced density matrix (RDM) to the spin sector, measuring quantum correlations due to the spin-orbit coupling (SOC). The analysis of IPR and entanglement entropies in terms of spin, wave number $k_{x}$ and position $y$, evidences a spin polarization structure and spatial confinement of near-gap wave functions at the boundaries $y=0,L$ and low $k_{x}$, as correspond to helical edge states. IPR localization measures provide momentum $k_{x}$ cutoffs from which near-gap states are no longer localized at the boundaries of the sample and become part of the bulk. Below this $k_{x}$-point cutoff, the entanglement entropy and the spin probabilities of the RDM also capture the spin polarization structure of edge states and exhibit a higher variability compared to the relatively low entropy of the bulk state region.
@@ I Introduction
In addition to the standard tight-binding method (see e.g. [3, 4, 5, 6, 7]), there are other analytic procedures in the current literature to study these finite size effects on edge states in the TI phase. Quadratic Hamiltonian expansions around $\Gamma$ or Dirac valley points $K$ in the Brillouin zone are used in [8, 6, 9] to analytically determine the expressions of the edge-state wavefunctions in a finite strip geometry $(x,y)\in(-\infty,\infty)\times[-L/2,L/2]$. Replacing the wavevector component $k_{y}$ by a complex number $-\mathrm{i}\lambda$, edge states display an exponential behavior $\psi(y)\sim\exp(\lambda y)$ to take into account an exponential localization at the boundaries $y=\pm L/2$. When the coherence length $1/\mathrm{Re}(\lambda)$ of edge states is comparable to the strip width $L$, states from opposite edges can overlap, leading to a gap opening in the edge state spectrum. These near-gap states exhibit localization properties typical of edge-states for low $k_{x}$ but they become degenerate with the bulk states from a certain $k$-point cutoff onwards where $\lambda$ becomes imaginary. However, this analysis is local (only valid in a small neighborhood around $\Gamma$ or $K$ points) and has some limitations.
@@
In this paper, we tackle the problem of finite size effects in the HgTe/CdTe semiconductors from a different (quantum information) perspective, including SOC due to bulk- and structure-inversion asymmetries (resp. BIA and SIA). The introduction of SOC has also been investigated in [7], where they use the tight-binding method to determine and exponential decay of the edge energy gap (with oscillations/modulations) with the strip width $L$ and to prove that this gap is not exactly localized at the $\Gamma$ point of the first Brillouin zone. This energy gap is also affected by an external perpendicular electric field, which tunes the Rashba (SIA) term of the Hamiltonian model and has nontrivial consequences on the charge conductance. We confirm this behavior for a more general BIA term including extra electron and hole couplings preserving time reversal symmetry. We also pursue the identification of topological order through quantum information (QI) measures and concepts like entanglement entropy. QI tools have also played an important role in the general understanding of quantum phase transitions. Indeed, entanglement is at the heart of the interplay between quantum information and quantum phases of matter (see e.g., [10, 11]). Signatures of topological phase transitions in higher Landau levels of HgTe/CdTe quantum wells without SOC from an information theory perspective have been reported in [12]. Other localization measures, like the inverse participation measure (IPR), has given useful information about the topological phase transition 2D Dirac materials like silicene [13]. This paper analyzes the structure of near-gap states in HgTe QWs with SOC under QI concepts like IPR and entanglement entropy, which turn out to be an interesting “microscope” to reveal details of their internal structure, and proposes a criterion for the differentiation between edge and bulk states near the gap attending to their IPR.
@@
The organization of the paper is as follows. In Sec. II we briefly discuss the structure of different HgTe QW Hamiltonian models considered in the literature and their topological phases. In Sec. III we approach the analysis of edge states in a finite strip geometry of width $L$ from two different perspectives: either looking for analytic localized eigenvectors of the low energy Hamiltonian, or by numerically solving the tight-binding model after a lattice regularization, comparing both approaches for some of the mentioned models. The first approach gives us a deeper understanding of the qualitative and internal structure of edge states (specially for the easier uncoupled case), but we shall rather follow the second approach to extract quantitative information, firstly about the spectrum and the dependence of the energy gap on the strip width $L$ and the Rashba coupling $\xi$, and its non-trivial consequences on the charge conductance of edge states and its potential use in the design of a QSH field effect transistor. In Sec. IV we take a closer look to the localization properties of near-gap states as a function of the spin ($s=\pm 1$), the momentum wave vector $k_{x}\in[-\pi/a,\pi/a]$, with $a$ the lattice constant, and the position $y\in[0,L]$ between the strip boundaries $y=0,L$. This study sheds light on the spin-polarization structure of edge states at the boundaries. The spreading of near-gap states in momentum ($k_{x}$) and position ($y$) space is analyzed through the IPR concept of QI, which allows to propose a criterion for the differentiation between edge-like and bulk-like behavior, attending to their localization properties. Finally, in Sec. V we use the reduced density matrix (RDM) to the spin sector to analyze spin up/down and spin-transfer probability densities of near-gap states as a function of the momentum $k_{x}$ and position $y$, paying especial attention to extremal values. This analysis also sheds light on the spin-polarization structure of edge states. The purity of the RDM (or equivalently, the linear entropy) also gives us information about the degree of entanglement between spin and band (electron-hole) sectors. Extremal entanglement values occur for special values of the position $y$ and momentum $k_{x}$. Other alternative correlation measures are also analyzed, all of them giving equivalent results. Finally, Sec. VI is devoted to conclusions.
@@ II.1 The uncoupled case
$H_{\mathrm{BHZ}}$ $=$ $\frac{\sigma_{0}+\sigma_{z}}{2}\otimes h_{+1}+\frac{\sigma_{0}-\sigma_{z}}{2}\otimes h_{-1},$
$h_{s}(\bm{k})$ $=$ $\epsilon_{0}(\bm{k})\sigma_{0}+\bm{d}_{s}(\bm{k})\cdot\bm{\sigma},\,s=\pm 1,$ (1)
where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ are Pauli matrices together with the $2\times 2$ identity matrix $\sigma_{0}$ and $\bm{k}=(k_{x},k_{y})$ is the wavevector. The spin $s=\pm 1$, $2\times 2$ matrix Hamiltonians $h_{s}(\bm{k})$ are related by $h_{-1}(\bm{k})=h_{+1}^{*}(-\bm{k})$ (temporarily reversed) and they admit an expansion around the center $\Gamma$ of the first Brillouin zone (FBZ) given by [18],
@@
where (+)(+)( + ) makes reference to conduction/electron and ()(-)( - ) to valence/hole states. The corresponding Hamiltonian $h_{s}$ spinor eigenvectors are
@@
$\Theta=-i(\sigma_{y}\otimes\sigma_{0})K,$ (5)
@@ II.2 Spin-orbit coupling
$H_{\mathrm{BIA}}(\bm{k})$ $=$ $\Delta_{z}(\sigma_{y}\otimes\sigma_{y})$
$+\frac{\Delta_{e}}{2}(k_{x}\sigma_{x}-k_{y}\sigma_{y})\otimes(\sigma_{0}+\sigma_{z})$
$+\frac{\Delta_{h}}{2}(k_{x}\sigma_{x}+k_{y}\sigma_{y})\otimes(\sigma_{0}-\sigma_{z}),$
$H_{\mathrm{SIA}}(\bm{k})$ $=$ $\frac{\xi}{2}(k_{x}\sigma_{y}-k_{y}\sigma_{x})\otimes(\sigma_{0}+\sigma_{z}).$ (9)
@@
where $\uparrow$,$\downarrow$ , ↓ makes reference to the spin degree of freedom $s=\pm 1$ and $EH$ denotes the electron and hole bands, respectively.
@@ III.1 Analytic approach to the solution of the effective continuous 4-band model
for the $n=2$ uncoupled case with $2\times 2$ Hamiltonian $h_{s}$ in (II.1), which yields $42n=$ 4 roots $\lambda_{1}=\lambda_{+},\lambda_{2}=\lambda_{-},\lambda_{3}=-\lambda_{+},\lambda_{4}=-\lambda_{-}$.
@@
For the material parameters in Table 1, Figure 1 shows the energy $E(k)$ of the edge states (in black color) of the uncoupled $n=2$ BHZ case for two values, $L_{1}=120$ nm (dashed) and $L_{2}=500$ nm (solid), of the strip width. The bulk conduction band for $h_{s}$ eigenstates are plotted in red for several values of $k_{y}$, highlighting the case $k_{y}=0=\lambda$ in magenta, where edge and bulk eigenstates merge at the cutoff wave numbers ±$\pm k_{1}$ (for strip width $L_{1}$) and ±$\pm k_{2}$ (for strip width $L_{2}$), marked by round black dots. Edge states show a noticeable energy gap for small $L$ ($E_{g}\simeq 2.27$ meV for $L=120$ nm) and a linear energy-momentum dispersion relation for large $L$, for which the gap tends to zero (see next paragraph). In Figure 2 we plot the cutoff wave numbers $k$ for several values of the strip width $L\in[100,900]$ nm, together with a rational fitting of type
@@
Figure 1: Energy of edge (black) and bulk (red for $k_{y}\not=0$ and magenta for $k_{y}=0$) Hamiltonian $h_{s}$ eigenstates as a function of $k_{x}$. Edge and bulk eigenstates merge at the cutoff wave numbers ±$\pm k_{1}=\pm 0.0096$ nm-1 (for strip width $L_{1}=120$nm, dashed black) and ±$\pm k_{2}=\pm 0.021$ nm-1 (for strip width $L_{2}=500$ nm, solid black), marked by round black dots. The corresponding energies are $E_{1}=10.58$ meV and $E_{2}=12.59$ meV. As a comparison, results from the numerical diagonalization of tight-binding model of Sec. III.2 are also plotted as black triangles for $L=120$ nm and black squares for $L=500$ nm.
@@
Figure 3: Density distribution $|\psi_{\uparrow E}|^{2}+|\psi_{\uparrow H}|^{2}$ of the first (conduction) spin up ($s=1$) edge state of the BHZ Hamiltonian $h_{s}$ for a strip width of $L=120$ nm as a function of $y$. The solid line corresponds to $k_{x}=0$ nm-1 ($E(0)\simeq 8.88$ meV), the dotted line to $k=0.005$ nm-1 and the dashed line to $k=-0.005$ nm-1 ($E(\pm 0.005)\simeq 9.48$ meV).
The dominant $\lambda$ root is the one with the larger length scale $|\mathrm{Re}(\lambda)|^{-1}$. The density distribution at the two edges $y=\pm L/2$ is mainly determined by the dominant root. We observe that spin up conduction states are localized at the boundary $y=L/2$ (resp. $y=-L/2$) for negative (resp. positive) $k_{x}$, the case $k_{x}=0$ displaying a more balanced configuration. This is a reflect of the already commented spin polarization structure of edge states (see later in Fig. 10 for more details).
@@ III.2 Lattice regularization and numerical diagonalization of the tight-binding model
[…]
@@
in position (discrete) $y$ and momentum $k=k_{x}$ spaces. Here we are considering a space discretization of the finite strip with $y_{n}=na,n=0,\dots,N=L/a$. The $4\times 4$ matrix $\mathcal{E}(k)$ results from eliminating all terms depending on $k_{y}$ in the regularized total Hamiltonian $H(\bm{k})$ in (10). Those terms then contribute to the matrix
@@
Notice that $\Psi_{1}$ and $\Psi_{2}$ are nearly degenerated for conduction and valence bands, but the energy $E_{1c}$ is a bit lower than $E_{2c}$ and $E_{1v}$ is slightly higher than $E_{2c}$, so that the energy gap is determined by $E_{g}=\min_{k}[E_{1c}(k)-E_{1v}(k)]$, with $k\in(-\pi/a,\pi/a)$. As we already anticipated in Sec. III.1, this gap shows an exponential decay with modulations/oscillations as function of the strip width $L$, as showed in Figure 5 (red dots). The value of $k$ minimizing $[[E_{1c}(k)-E_{1v}(k)]$ does not coincide in general with $k=0$ (the $\Gamma$ point) but it is quite close to it. To compute it, one needs a dense mesh of points $k_{n}$ which requires high computational resources for high $L$. Thankfully, larger values of the SOC coupling $\Delta_{z}$ require smaller Hamiltonian matrix sizes $N=L/a$ and less computational effort to observe energy gap oscillations. We have chosen $\Delta_{z}=10$ meV this time (different from the $\Delta_{z}=1.6$ meV value in Table 1 used in Figure 4) just for computational convenience. Sudden gap drops occur at the critical strip widths $L_{c}\simeq 100$ and $L_{c}\simeq 220$nm. The exponential decay of the energy gap is captured by the gap $E_{g}^{\Gamma}=[E_{1c}(0)-E_{1v}(0)]$ at the $\Gamma$ point (black dots in Figure 5). A fit of nine values of $E_{g}^{\Gamma}$ at $L=100,\dots,500$, in steps of $\Delta L=50$, provides the expression

### v3

@@ Abstract
Quantum information measures are proposed to analyze the structure of near-gap electronic states in HgTe quantum wells in a strip geometry $(x,y)\in(-\infty,\infty)\times[0,L]$ of finite width $L$. This allows us to establish criteria for distinguishing edge from bulk states in the topological insulator phase, including the transition region and cutoff of the wave number $k_{x}$ where edge states degenerate with bulk states. Qualitative and quantitative information on the near-gap Hamiltonian eigenstates, obtained by tight-binding calculations, is extracted from localization measures, like the inverse participation ratio (IPR), entanglement entropies of the reduced density matrix (RDM) to the spin sector –measuring quantum correlations due to the spin-orbit coupling (SOC)– and from correlation functions for a $y$-space partition. The analysis of IPR and entanglement entropies in terms of spin, wave number $k_{x}$ and position $y$, evidences a spin polarization structure and spatial confinement of near-gap wave functions at the boundaries $y=0,L$ and low $k_{x}$, as correspond to helical edge states. IPR localization measures provide momentum $k_{x}$ cutoffs from which near-gap states are no longer localized at the boundaries of the sample and become part of the bulk. Below this $k_{x}$-point cutoff, the entanglement entropy and the spin probabilities of the RDM also capture the spin polarization structure of edge states and exhibit a higher variability compared to the relatively low entropy of the bulk state region. For a real-space partition, the edge-state region in momentum space exhibits lower correlation modulus, but higher correlation arguments, than the bulk-state region.
edge states, topological insulators, spin-orbit coupling, quantum information, reduced density matrix, correlations.
@@ I Introduction
In addition to the standard tight-binding method (see e.g. [3, 4, 5, 6, 7]), there are other analytic procedures in the current literature to study these finite size effects on edge states in the TI phase. Quadratic Hamiltonian expansions around $\Gamma$ or Dirac valley points $K$ and $K^{\prime}$ in the Brillouin zone are used in [8, 6, 9] to analytically determine the expressions of the edge-state wavefunctions in a finite strip geometry. For the analytic approach in Section III.1, the symmetric choice $y\in[-L/2,L/2]$ seems to be more (computationally) convenient than $y\in[0,L]$, although both conventions are equivalent. Replacing the wavevector component $k_{y}$ by a complex number $\mathrm{i}\lambda$, edge states display an exponential behavior $\psi(y)\sim\exp(\lambda y)$ to take into account an exponential localization at the boundaries $y=\pm L/2$. When the coherence length $1/\mathrm{Re}(\lambda)$ of edge states is comparable to the strip width $L$, states from opposite edges can overlap, leading to a gap opening in the edge state spectrum. These near-gap states exhibit localization properties typical of edge-states for low $k_{x}$ but they become degenerate with the bulk states from a certain $k$-point cutoff onwards where $\lambda$ becomes imaginary. However, this analysis is local (only valid in a small neighborhood around $\Gamma$ or $K$ points) and has some limitations.
@@
In this paper, we tackle the problem of finite size effects in the HgTe/CdTe semiconductors from a different (quantum information) perspective, including SOC due to bulk- and structure-inversion asymmetries (resp. BIA and SIA). The introduction of SOC has also been investigated in [7], where they use the tight-binding method to determine an exponential decay of the edge energy gap (with oscillations/modulations) with the strip width $L$ and to prove that this gap is not exactly localized at the $\Gamma$ point of the first Brillouin zone. This energy gap is also affected by an external perpendicular electric field, which tunes the Rashba (SIA) term of the Hamiltonian model and has nontrivial consequences on the charge conductance. We confirm this behavior for a more general BIA term including extra electron and hole couplings preserving time reversal symmetry. We also pursue the identification of topological order through quantum information (QI) measures and concepts like entanglement entropy. QI tools have also played an important role in the general understanding of quantum phase transitions. Indeed, entanglement is at the heart of the interplay between quantum information and quantum phases of matter (see e.g., [10, 11]). Some standard literature on entanglement research for topological phases primarily examines real-space partitions, revealing universal information of topological phases [12, 13]. The entanglement energy spectrum analysis in [14] unveils edge state information through momentum-resolved counting of entanglement energy levels, as captured by the “edge-entanglement spectrum correspondence” [15] or the Li-Haldane conjecture [16]. Signatures of topological phase transitions in higher Landau levels of HgTe/CdTe quantum wells without SOC from an information theory perspective have been reported in [17]. Other localization measures, like the inverse participation measure (IPR), has given useful information about the topological phase transition 2D Dirac materials like silicene [18]. In this paper we adopt a less standard spin-orbital bipartition to analyze the structure of near-gap states in HgTe QWs with SOC under QI concepts like IPR and entanglement entropy, which turn out to be an interesting “microscope” to reveal details of their internal structure, and proposes a criterion for the differentiation between edge and bulk states near the gap attending to their IPR.
@@
The organization of the paper is as follows. In Sec. II we briefly discuss the structure of different HgTe QW Hamiltonian models considered in the literature and their topological phases. In Sec. III we approach the analysis of edge states in a finite strip geometry of width $L$ from two different perspectives: either looking for analytic localized eigenvectors of the low energy Hamiltonian, or by numerically solving the tight-binding model after a lattice regularization, comparing both approaches for some of the mentioned models. The first approach gives us a deeper understanding of the qualitative and internal structure of edge states (specially for the easier uncoupled case), but we shall rather follow the second approach to extract quantitative information, firstly about the spectrum and the dependence of the energy gap on the strip width $L$ and the Rashba coupling $\xi$, and its non-trivial consequences on the charge conductance of edge states and its potential use in the design of a QSH field effect transistor. In Sec. IV we take a closer look to the localization properties of near-gap states as a function of the spin ($s=\pm 1$), the momentum wave vector $k_{x}\in[-\pi/a,\pi/a]$, with $a$ the lattice constant, and the position $y\in[0,L]$ between the strip boundaries $y=0,L$. This study sheds light on the spin-polarization structure of edge states at the boundaries. The spreading of near-gap states in momentum ($k_{x}$) and position ($y$) space is analyzed through the IPR concept of QI, which allows to propose a criterion for the differentiation between edge-like and bulk-like behavior, attending to their localization properties. Finally, in Sec. V we use the reduced density matrix (RDM) to the spin sector to analyze spin up/down and spin-transfer probability densities of near-gap states as a function of the momentum $k_{x}$ and position $y$, paying especial attention to extremal values. This analysis also sheds light on the spin-polarization structure of edge states. The purity of the RDM (or equivalently, the linear entropy) also gives us information about the degree of entanglement between spin and band (electron-hole) sectors. Extremal entanglement values occur for special values of the position $y$ and momentum $k_{x}$. Other alternative correlation measures are also analyzed, all of them giving equivalent results. Section VI is devoted to correlation functions for a real-space partition. Finally, Sec. VII is devoted to conclusions.
@@ II.1 The uncoupled case
$H_{\mathrm{BHZ}}$ $=$ $\frac{\sigma_{0}+\sigma_{z}}{2}\otimes h_{+1}+\frac{\sigma_{0}-\sigma_{z}}{2}\otimes h_{-1}=\begin{pmatrix}h_{+1}&0\\ 0&h_{-1}\end{pmatrix}$,
$h_{s}(\bm{k})$ $=$ $\epsilon_{0}(\bm{k})\tau_{0}+\bm{d}_{s}(\bm{k})\cdot\bm{\tau},\;s=\pm 1,$ (1)
where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$, together with the $2\times 2$ identity matrix $\sigma_{0}$, are Pauli matrices acting on the spin basis, and $\bm{\tau}=(\tau_{x},\tau_{y},\tau_{z})$ acting on the electron-hole subbands. The wavevector is denoted by $\bm{k}=(k_{x},k_{y})$. The spin $s=\pm 1$, $2\times 2$ matrix Hamiltonians $h_{s}(\bm{k})$ are related by $h_{-1}(\bm{k})=h_{+1}^{*}(-\bm{k})$ (temporarily reversed) and they admit an expansion around the center $\Gamma$ of the first Brillouin zone (FBZ) given by [23],
@@
where $(+)$ makes reference to conduction/electron and ()(-) to valence/hole states. The corresponding Hamiltonian $h_{s}$ spinor eigenvectors are
@@
$\Theta=-i(\sigma_{y}\otimes\tau_{0})K,$ (5)
@@ II.2 Spin-orbit coupling
$H_{\mathrm{BIA}}(\bm{k})$ $=$ $\Delta_{z}(\sigma_{y}\otimes\tau_{y})$
$+\frac{\Delta_{e}}{2}(k_{x}\sigma_{x}-k_{y}\sigma_{y})\otimes(\tau_{0}+\tau_{z})$
$+\frac{\Delta_{h}}{2}(k_{x}\sigma_{x}+k_{y}\sigma_{y})\otimes(\tau_{0}-\tau_{z}),$
$H_{\mathrm{SIA}}(\bm{k})$ $=$ $\frac{\xi}{2}(k_{x}\sigma_{y}-k_{y}\sigma_{x})\otimes(\tau_{0}+\tau_{z}).$ (9)
@@
where $\uparrow,\downarrow$ makes reference to the spin degree of freedom $s=\pm 1$ and $EH$ denotes the electron and hole bands, respectively.
@@ III.1 Analytic approach to the solution of the effective continuous 4-band model
for the $n=2$ uncoupled case with $2\times 2$ Hamiltonian $h_{s}$ in (II.1), which yields $2n=4$ roots $\lambda_{1}=\lambda_{+},\lambda_{2}=\lambda_{-},\lambda_{3}=-\lambda_{+},\lambda_{4}=-\lambda_{-}$.
@@
For the material parameters in Table 1, Figure 1 shows the energy $E(k)$ of the edge states (in black color) of the uncoupled $n=2$ BHZ case for two values, $L_{1}=120$ nm (dashed) and $L_{2}=500$ nm (solid), of the strip width. The bulk conduction band for $h_{s}$ eigenstates are plotted in red for several values of $k_{y}$, highlighting the case $k_{y}=0=\lambda$ in magenta, where edge and bulk eigenstates merge at the cutoff wave numbers $\pm k_{1}$ (for strip width $L_{1}$) and $\pm k_{2}$ (for strip width $L_{2}$), marked by round black dots. Edge states show a noticeable energy gap for small $L$ ($E_{g}\simeq 2.27$ meV for $L=120$ nm) and a linear energy-momentum dispersion relation for large $L$, for which the gap tends to zero (see next paragraph). In Figure 2 we plot the cutoff wave numbers $k$ for several values of the strip width $L\in[100,900]$ nm, together with a rational fitting of type
@@
Figure 1: Energy of edge (black) and bulk (red for $k_{y}\not=0$ and magenta for $k_{y}=0$) Hamiltonian $h_{s}$ eigenstates as a function of $k_{x}$. Edge and bulk eigenstates merge at the cutoff wave numbers $\pm k_{1}=\pm 0.0096$ nm-1 (for strip width $L_{1}=120$nm, dashed black) and $\pm k_{2}=\pm 0.021$ nm-1 (for strip width $L_{2}=500$ nm, solid black), marked by round black dots. The corresponding energies are $E_{1}=10.58$ meV and $E_{2}=12.59$ meV. As a comparison, results from the numerical diagonalization of tight-binding model of Sec. III.2 are also plotted as black triangles for $L=120$ nm and black squares for $L=500$ nm.
@@
Figure 3: Density distribution $|\psi_{\uparrow E}|^{2}+|\psi_{\uparrow H}|^{2}$ of the first (conduction) spin up ($s=1$) edge state of the BHZ Hamiltonian $h_{s}$ for a strip width of $L=120$ nm as a function of $y$. The solid line corresponds to $k_{x}=0$ nm-1 ($E(0)\simeq 8.88$ meV), the dotted line to $k=0.005$ nm-1 and the dashed line to k=0.005k=-0.005 nm-1 ($E(\pm 0.005)\simeq 9.48$ meV).
The dominant $\lambda$ root is the one with the larger length scale $|\mathrm{Re}(\lambda)|^{-1}$. The density distribution at the two edges $y=\pm L/2$ is mainly determined by the dominant root. We observe that spin up conduction states are localized at the boundary $y=L/2$ (resp. y=L/2y=-L/2) for negative (resp. positive) $k_{x}$, the case $k_{x}=0$ displaying a more balanced configuration. This is a reflect of the already commented spin polarization structure of edge states (see later in Fig. 10 for more details).
@@ III.2 Lattice regularization and numerical diagonalization of the tight-binding model
in the Hamiltonian $H(\bm{k})$ in (10), with $a$ the lattice constant. Then, the Brillouin zone (BZ) is $\bm{k}\in(-\pi/a,\pi/a)\times(-\pi/a,\pi/a)$. We shall set $a=2$ nm in all our calculations, in order to compare some of our results with those of Ref. [7], whose authors also use this value. Therefore, we shall consider a space discretization of the finite strip along the $y$-direction, with $y_{n}=na$ and $n=0,\dots,N=L/a$ the site indices, and we shall place hereinafter the lattice boundaries at $y=0,L$, accordingly. The translation between nanometers and lattice units can then be done by simply dividing lengths by $a=2$.
Following the general procedure of Refs. [6, 7], one Fourier transforms $k_{y}$ in the total Hamiltonian $\mathcal{H}=\int_{\mathrm{BZ}}d\bm{k}H(\bm{k})c^{\dagger}_{\bm{k}}c_{\bm{k}}$ by substituting the annihilation (viz. creation) operators
@@
in position (discrete) $y$ and momentum $k=k_{x}$ spaces. The $4\times 4$ matrix $\mathcal{E}(k)$ results from eliminating all terms depending on $k_{y}$ in the regularized total Hamiltonian $H(\bm{k})$ in (10). Those terms then contribute to the matrix
@@
Notice that $\Psi_{1}$ and $\Psi_{2}$ are nearly degenerated for conduction and valence bands, but the energy $E_{1c}$ is a bit lower than $E_{2c}$ and $E_{1v}$ is slightly higher than $E_{2c}$, so that the energy gap is determined by $E_{g}=\min_{k}[E_{1c}(k)-E_{1v}(k)]$, with $k\in(-\pi/a,\pi/a)$. As we already anticipated in Sec. III.1, this gap shows an exponential decay with modulations/oscillations as function of the strip width $L$, as showed in Figure 5 (red dots). The value of $k$ minimizing $E_{1c}(k)-E_{1v}(k)]$ does not coincide in general with $k=0$ (the $\Gamma$ point) but it is quite close to it. To compute it, one needs a dense mesh of points $k_{n}$ which requires high computational resources for high $L$. Thankfully, larger values of the SOC coupling $\Delta_{z}$ require smaller Hamiltonian matrix sizes $N=L/a$ and less computational effort to observe energy gap oscillations. We have chosen $\Delta_{z}=10$ meV this time (different from the $\Delta_{z}=1.6$ meV value in Table 1 used in Figure 4) just for computational convenience. Sudden gap drops occur at the critical strip widths $L_{c}\simeq 100$ and $L_{c}\simeq 220$nm. The exponential decay of the energy gap is captured by the gap $E_{g}^{\Gamma}=[E_{1c}(0)-E_{1v}(0)]$ at the $\Gamma$ point (black dots in Figure 5). A fit of nine values of $E_{g}^{\Gamma}$ at $L=100,\dots,500$, in steps of $\Delta L=50$, provides the expression